OpenStudy (anonymous):

differentiate using implicit diff tan^3 (xy^2+y)=x

OpenStudy (dumbcow):

is this same question?

OpenStudy (anonymous):

yea i forgot the brackets!

OpenStudy (dumbcow):

oh so my solution is completly wrong then, haha :)

OpenStudy (anonymous):

pliz try to solve it again

OpenStudy (dumbcow):

you have to use chain rule \[(xy^{2} +y)' = y^{2} + 2xy \frac{dy}{dx} + \frac{dy}{dx}\] \[\rightarrow 3\tan^{2}(xy^{2}+y) \sec^{2}(xy^{2}+y)(y^{2} + 2xy \frac{dy}{dx} + \frac{dy}{dx}) = 1\] For simplicity, let u = tan^2*sec^2 \[\rightarrow 3uy^{2} +6uxy \frac{dy}{dx}+3u \frac{dy}{dx} = 1\] \[\frac{dy}{dx} = \frac{1-3uy^{2}}{3u(2xy+1)}\]